Five hundred children participated in a field demonstration. Their heights averaged is 110 cm with a standard deviation of 6 cm. How many children have heights between 98 cm and 104 cm?


1
Expert's answer
2021-04-29T18:00:16-0400

mean "= \mu = 110"

standard deviation "= \sigma = 6"

"P(98<x<104) = P(\frac{98-110}{6}<Z< \frac{104-110}{6}) \\

= P(-2 <Z< -1) \\

= P(Z< -1) -P(Z< -2) \\

= 0.1586 – 0.0227 \\

= 0.1359"

= 13.59 %

"500 \times 0.1359 = 68" children have heights between 98 cm and 104 cm